# Towards Understanding the Mechanisms of Classifier-Free Guidance

### v1

@@ Abstract
Classifier-free guidance (CFG) is a core technique powering state-of-the-art image generation systems, yet its underlying mechanisms remain poorly understood. In this work, we begin by analyzing CFG in a simplified linear diffusion model, where we show its behavior closely resembles that observed in the nonlinear case. Our analysis reveals that linear CFG improves generation quality via three distinct components: (i) a mean-shift term that approximately steers samples in the direction of class means, (ii) a positive Contrastive Principal Components (CPC) term that amplifies class-specific features, and (iii) a negative CPC term that suppresses generic features prevalent in unconditional data. We then verify that these insights in real-world, nonlinear diffusion models: over a broad range of noise levels, linear CFG resembles the behavior of its nonlinear counterpart. Although the two eventually diverge at low noise levels, we discuss how the insights from the linear analysis still shed light on the CFG’s mechanism in the nonlinear regime.
@@ 1 Introduction
$\mathbb{E}_{\bm{x}\sim p_{\text{data}},\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}(t)\bm{I})}[||\mathcal{D}_{\bm{\theta}}(\bm{x}+\bm{\epsilon};\sigma(t))-\bm{x}||_{2}^{2}].$ (3)
@@ 2.1 Optimal Linear Diffusion Model
where $\bm{\Sigma}=\bm{U}\bm{\Lambda}\bm{U}^{T}$ is the full SVD of the covariance matrix, $\bm{\Lambda}=\text{diag}(\lambda_{1},\cdots,\lambda_{d})$ is the singular values and $\tilde{\bm{\Lambda}}_{\sigma(t)}=\operatorname{diag}\left(\frac{\lambda_{1}}{\lambda_{1}+\sigma^{2}(t)},\cdots,\frac{\lambda_{d}}{\lambda_{d}+\sigma^{2}(t)}\right)$. With this linear denoiser, the reverse diffusion ODE (2) has the following closed-form expression (see section B.1 for the proof):
@@
As shown in Figures 1 and 15, the sampling trajectories of the deep diffusion model (EDM) and the linear model share high similarity at high to moderate noise levels. Although the models begin to diverge at lower noise levels—where EDM exhibits strong nonlinearity and realistic image content begins to form—their final samples still share a similar overall structure. Moreover, as shown in [21], this similarity is particularly obvious when the deep network has limited capacity or the training is insufficient. Since $\mathcal{D}_{\mathrm{L}}(\bm{x};\sigma(t))$ is the optimal denoiser for $p(\bm{x};\sigma(t))$ induced by $p_{\text{data}}(\bm{x})=\mathcal{N}(\bm{\mu},\bm{\Sigma})$, sampling with $\mathcal{D}_{\mathrm{L}}$ is equivalent to sampling from $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. Hence, we refer to $\mathcal{D}_{\mathrm{L}}$ as the linear Gaussian model.
@@ 2.2 Contrastive Principal Component Analysis
which are essentially the top eigenvectors of $\bm{\Sigma}_{X}-\bm{\Sigma}_{Y}$. Geometrically, the first $k$ CPCs span the $k$-dimensional subspace that best fits the dataset $X$ while being as far as possible from $Y$ (see appendix A for details). Conversely, directions $\bm{v}$ for which $\text{Var}_{X}(\bm{v})\approx\text{Var}_{Y}(\bm{v})$ represent either universal structures shared by both $X$ and $Y$ or meaningless features lying in the null space of the data covariances—and are thus discarded as less interesting. Finally, a scalar factor can be introduced in (10) to control the strength of the contrast.
@@ 3 Analyzing CFG in Linear Model
Figure 2: Effects of CFG. Left and right compare naive conditional sampling (top rows) versus CFG-guided sampling (bottom rows) for deep diffusion models (EDM) and linear Gaussian diffusion models, respectively. Each grid ceil corresponds to the same initial noise. While naive conditional samples lack class-specific clarity, CFG significantly improves both visual quality and distinctiveness. The conditional linear models are built with class-specific means and covariances. Please refer to appendix D for more experiment results.
In this section, we first show that naive conditional sampling often produces low-quality samples lacking clear class-specific features, which we attribute to the non-distinctiveness of class covariance matrices (section 3.1). We then theoretically analyze how CFG in the context of linear diffusion models alleviates this issue (section 3.2).
@@ 3.1 Naive Conditional Generation Lacks Class-Specificity
Figure 2(left) (top row) shows the samples generated via naive conditional sampling (4). Qualitatively, these samples often exhibit poor image quality, with incoherent features that blend into the background and the class-specific image structures can be hard to discriminate. Moreover, even when conditioned on different class labels, images generated from the same initial noise share high structural similarity. This indicates that naive conditional sampling fails to capture distinctive, class-specific patterns.
@@
To quantify this class-specificity gap, we compare the pairwise class similarity with the FID score [27]. For a pair of two classes, we build two datasets $X$ and $Y$ and compute the FID between them. As shown in Figure 3, when $X$ and $Y$ are built with images generated with naive conditional sampling, the FID (colored in orange) is consistently lower than when they are built with the training data (colored in blue). Because lower FID indicates higher similarity, this result confirms that compared with the training images, which represent the ground truth data distribution, images generated by naive conditional sampling are less distinguishable across classes.
@@
Figure 3: Class-to-Class Similarity. Each cell reports the FID between datasets of two classes, built with (i) training data (ii) data generated by naive conditional sampling and (iii) data generated by CFG sampling (refer to section D.2 for experiment details and more results.)
@@
This issue is particularly evident in linear diffusion models. As shown in Figure 2(right, top row), samples generated with linear diffusion models built with class-specific means and covariances look highly similar. From (8), we see that the linear sampling trajectory is governed by the data covariance: $\bm{x}_{t}$ is a linear combination of PCs, weighted by (i) the correlation $\bm{u}_{i}^{T}(\bm{x}_{T}-\bm{\mu})$ between the mean-subtracted initial noise $\bm{x}_{T}$ and the $i$-th PC, and (ii) scaling factors $\sqrt{\frac{\lambda_{i}+\sigma^{2}(t)}{\lambda_{i}+\sigma^{2}(T)}}$ that emphasize leading PCs. Consequently, if the class-conditional covariances lack sufficiently discriminative structures (which is indeed true as shown in section D.3), generated samples will appear similar regardless of class label. This lack of class-specificity aligns with prior findings [25], which shows that PCs often capture generic image variations (e.g., foreground-background), rather than class-specific patterns.
@@
We hypothesize that real-world diffusion models inherit similar limitations. As discussed in section 2.1, for high to moderate noise levels, deep diffusion models can be well approximated by linear models, especially when network has limited capacity and training is insufficient. Figure 2 (top row) and Figure 14, 15 demonstrate that linear models reproduce the coarse-grained structures of nonlinear diffusion samples, implying that the covariance structure plays a significant role in shaping the high-level features of the generated samples. Hence, if covariances are indistinct across classes, sample quality can be restricted even in nonlinear models (see section D.3 for further discussion).
@@
As quantitatively shown in Figure 3, CFG significantly increases the inter-class separation: FID (colored in green) between different generated classes rises. Qualitatively, Figure 2 (bottom row) shows that CFG substantially improves both linear and nonlinear models, producing visibly better samples with enhanced class-specific structures. Similar effects of CFG across both linear and nonlinear models motivate us to use a linear model as a simplified prototype to analyze how CFG reshapes the generation process and why it is effective.
@@ 3.2 How Linear CFG Leads to Distinct Generations
Here, $f_{c}(\bm{x}_{t},t)$ is the standard conditional score, and $g_{cpc}(\bm{x}_{t},t)$ plus $g_{mean}(t)$ form the CFG guidance (derivation for the decomposition is provided in section B.2). Let $\bm{V}_{\sigma(t)}\hat{\bm{\Lambda}}_{\sigma(t)}\bm{V}_{\sigma(t)}^{T}$ be the eigen decomposition of $\tilde{\bm{\Sigma}}_{c,t}-\tilde{\bm{\Sigma}}_{uc,t}$, whose spectrum contains both positive and negative eigenvalues (see Figure 19), $g_{cpc}(\bm{x}_{t},t)$ can be split accordingly into positive and negative CPC components:
@@
where $\bm{V}_{\sigma(t),+}$ and $\bm{V}_{\sigma(t),-}$ contain eigenvectors $\bm{v}_{+,i}$ and $\bm{v}_{-.i}$ corresponding to positive and negative eigenvalues $\hat{\bm{\Lambda}}_{\sigma(t),+}$ and $\hat{\bm{\Lambda}}_{\sigma(t),-}$ respectively. As discussed in section 2.3, $\tilde{\bm{\Sigma}}_{c,t}$ and $\tilde{\bm{\Sigma}}_{uc,t}$ are up to a scaling factor $\sigma^{2}(t)$ equivalent to the conditional and unconditional posterior covariances of $p_{\text{data}}(\bm{x}|\bm{c})=\mathcal{N}(\bm{\mu}_{c},\bm{\Sigma}_{c})$ and $p_{\text{data}}(\bm{x})=\mathcal{N}(\bm{\mu}_{uc},\bm{\Sigma}_{uc})$. Hence, $\bm{V}_{\sigma(t)}$ are the CPCs which contrast between $X\sim p_{data}(\bm{x}|\bm{x}_{t},\bm{c})$ and $Y\sim p_{\text{data}}(\bm{x}|\bm{x}_{t})$. Specifically, $\bm{V}_{\sigma(t),+}$ captures directions of higher conditional variance (class-specific features), while $\bm{V}_{\sigma(t),-}$ captures directions of higher unconditional variance (features more prevalent in the unconditional data).
@@
Distinctive Effects of the CFG Components. Figure 4(a) shows that for both nonlinear (EDM) and linear models, CFG significantly enhances the characteristic pattern—a person holding a fish—of the "tench" class from ImageNet [28]. Next, we isolate the roles of each CFG term by selectively enabling only one at a time within the linear model. In the following discussion, we omit the negative sign in (12) since the ODE runs backward in time:
@@
  • The positive CPC term (13) projects $\bm{x}_{t}-\bm{\mu}_{c}$ onto the subspace spanned by the positive CPCs, i.e., the eigenvectors $\bm{v}_{+,i}$ associated with positive eigenvalues, with each component scaled by its eigenvalue $\hat{\lambda}_{+,i}$ and the guidance strength $\gamma$. Since $\hat{\lambda}_{+,i}\geq 0$, (13) is added to $\bm{x}_{t}$, i.e., the components of $\bm{x}_{t}-\bm{\mu}_{c}$ that align with the positive CPCs, which represent the class-specific features, are amplified.  Figure 4 (b) (second column) show the first 25 positive CPCs of $\bm{\Sigma}_{c}-\bm{\Sigma}_{uc}$ 111 Although $\bm{V}_{\sigma(t)}$ depends on $\sigma(t)$, over a wide range of noise levels (especially high ones), it remains close to the eigenvectors of $\bm{\Sigma}_{c}-\bm{\Sigma}_{uc}$. We provide its full evolution across time in Figure 20. and the resulting samples. Compared to the conditional PCs of the dataset, the positive CPCs better capture the unique patterns of the class, which emerge visibly in the generated images.
@@
  • Similarly, the negative CPC term (14) projects $\bm{x}_{t}-\bm{\mu}_{c}$ onto the negative CPC directions $\bm{v}_{-,i}$. Since $\hat{\lambda}_{-,i}<0$, these components are subtracted from $\bm{x}_{t}$, suppressing features associated more strongly with the unconditional data. Figure 4 (b) (third column) shows the first 25 negative CPCs and the resulting generations. Although visually less interpretable, these directions represent common but target-class-irrelevant features in the unconditional data. Suppressing them reduces background clutter and irrelevant content, making class-relevant structures more salient.
@@
  • In the context of linear diffusion model, it can be shown that (see proof in section B.2):
@@
    Figure 4(b), fourth column, shows that the mean-shift term enhances the structure of class mean in the generated samples. However, unlike the positive CPC term, $g_{mean}(t)$ is independent of $\bm{x}_{t}$, thus producing more homogeneous samples with reduced diversity.
Figure 4: Distinct effects of different CFG components. (a) CFG substantially enhances class-specific features (in both EDM and linear diffusion). (b) Top row: PCs, positive/negative CPCs, and $\bm{\mu}_{c}-\bm{\mu}_{uc}$. Bottom row: generated samples when each component is applied in isolation. (c) One-dimensional densities of generated samples after projection onto key directions. The left column corresponds to the linear diffusion model, whereas the right column corresponds to the EDM model. Top row: project onto leading positive CPC. Middle row: project onto negative CPC. Third row: project onto the mean-shift direction. Here we only plot the resulting histograms for the first positive and negative CPCs but the same patterns hold for subsequent CPCs. For experimental details and more results, please refer to section E.3.
@@
The proof is postponed to appendix B. Compared to the solution of the standard conditional sampling (8), the CFG guidance introduces the following two effects:
@@
In Section B.3, we provide an empirical validation of Theorem 1 using a 2D synthetic dataset that satisfies Assumption 1. Here, we further verify the CFG’s effects of enhancing (suppressing) CPC components and shifting samples towards the mean-shift direction in natural image dataset through the following experiment:
@@
The above experiments are conducted on both linear and nonlinear (EDM) diffusion models. The resulting univariate distributions are shown in Figure 4(c). Compared with naive conditional sampling, CFG shifts probability mass toward higher projection values along the positive-CPC and mean-shift directions, and toward lower values along the negative-CPC direction, indicating that the first two are amplified whereas the third is suppressed.

### v3

@@ Abstract
Classifier-free guidance (CFG) is a core technique powering state-of-the-art image generation systems, yet its underlying mechanisms remain poorly understood. In this work, we begin by analyzing CFG in a simplified linear diffusion model, where we show its behavior closely resembles that observed in the nonlinear case. Our analysis reveals that linear CFG improves generation quality via three distinct components: (i) a mean-shift term that approximately steers samples in the direction of class means, (ii) a positive Contrastive Principal Components (CPC) term that amplifies class-specific features, and (iii) a negative CPC term that suppresses generic features prevalent in unconditional data. We then verify these insights in real-world, nonlinear diffusion models: over a broad range of noise levels, linear CFG resembles the behavior of its nonlinear counterpart. Although the two eventually diverge at low noise levels, we discuss how the insights from the linear analysis still shed light on the CFG’s mechanism in the nonlinear regime.
@@ 1 Introduction
[…]
@@ 2.1 Optimal Linear Diffusion Model
where $\bm{\Sigma}=\bm{U}\bm{\Lambda}\bm{U}^{T}$ is the full SVD of the covariance matrix, $\bm{\Lambda}=\text{diag}(\lambda_{1},\cdots,\lambda_{d})$ is the singular values and $\tilde{\bm{\Lambda}}_{\sigma(t)}=\operatorname{diag}\left(\frac{\lambda_{1}}{\lambda_{1}+\sigma^{2}(t)},\cdots,\frac{\lambda_{d}}{\lambda_{d}+\sigma^{2}(t)}\right)$. With this linear denoiser, the reverse diffusion ODE (2) has the following closed-form expression (see section˜B.1 for the proof):
@@
As shown in Figures˜1 and 15, the sampling trajectories of the deep diffusion model (EDM) and the linear model share high similarity at high to moderate noise levels. Although the models begin to diverge at lower noise levels—where EDM exhibits strong nonlinearity and realistic image content begins to form—their final samples still share a similar overall structure. Moreover, as shown in [21], this similarity is particularly obvious when the deep network has limited capacity or the training is insufficient. Since $\mathcal{D}_{\mathrm{L}}(\bm{x};\sigma(t))$ is the optimal denoiser for $p(\bm{x};\sigma(t))$ induced by $p_{\text{data}}(\bm{x})=\mathcal{N}(\bm{\mu},\bm{\Sigma})$, sampling with $\mathcal{D}_{\mathrm{L}}$ is equivalent to sampling from $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. Hence, we refer to $\mathcal{D}_{\mathrm{L}}$ as the linear Gaussian model.
@@ 2.2 Contrastive Principal Component Analysis
which are essentially the top eigenvectors of $\bm{\Sigma}_{X}-\bm{\Sigma}_{Y}$. Geometrically, the first $k$ CPCs span the $k$-dimensional subspace that best fits the dataset $X$ while being as far as possible from $Y$ (see appendix˜A for details). Conversely, directions $\bm{v}$ for which $\text{Var}_{X}(\bm{v})\approx\text{Var}_{Y}(\bm{v})$ represent either universal structures shared by both $X$ and $Y$ or meaningless features lying in the null space of the data covariances—and are thus discarded as less interesting. Finally, a scalar factor can be introduced in (10) to control the strength of the contrast.
@@ 3 Analyzing CFG in Linear Model
Figure 2: Effects of CFG. Left and right compare naive conditional sampling (top rows) versus CFG-guided sampling (bottom rows) for deep diffusion models (EDM) and linear Gaussian diffusion models, respectively. Each grid ceil corresponds to the same initial noise. While naive conditional samples lack class-specific clarity, CFG significantly improves both visual quality and distinctiveness. The conditional linear models are built with class-specific means and covariances. Please refer to appendix˜D for more experiment results.
In this section, we first show that naive conditional sampling often produces low-quality samples lacking clear class-specific features, which we attribute to the non-distinctiveness of class covariance matrices (section˜3.1). We then theoretically analyze how CFG in the context of linear diffusion models alleviates this issue (section˜3.2).
@@ 3.1 Naive Conditional Generation Lacks Class-Specificity
Figure˜2(left) (top row) shows the samples generated via naive conditional sampling (4). Qualitatively, these samples often exhibit poor image quality, with incoherent features that blend into the background and the class-specific image structures can be hard to discriminate. Moreover, even when conditioned on different class labels, images generated from the same initial noise share high structural similarity, suggesting that naive conditional sampling fails to capture discriminative, class-dependent patterns.
@@
To quantify this loss of class-specificity, we compute the pairwise inter-class similarity with the FID metric [27]. For each pair of classes, we construct two datasets $X$ and $Y$ and evaluate the FID between them. As shown in Figure˜3, when $X$ and $Y$ are built with images generated with naive conditional sampling, the FID (colored in orange) is consistently lower than when they are built with the training data (colored in blue). Since lower FID indicates higher similarity, this result confirms that compared with the training images, which represent the ground truth data distribution, images generated by naive conditional sampling are less distinguishable across classes.
@@
Figure 3: Class‑to‑Class Similarity. Each cell reports the FID between datasets of two classes, built with (i) training data (ii) data generated by naive conditional sampling and (iii) data generated by CFG sampling (refer to section˜D.2 for experiment details and more results.)
@@
This issue is especially pronounced in linear diffusion models. As shown in Figure˜2(right, top row), samples generated with linear diffusion models built with class-specific means and covariances appear highly similar. From (8), we see that the linear sampling trajectory is governed by the data covariance: $\bm{x}_{t}$ is a linear combination of PCs, weighted by (i) the correlation $\bm{u}_{i}^{T}(\bm{x}_{T}-\bm{\mu})$ between the mean-subtracted initial noise $\bm{x}_{T}$ and the $i$-th PC, and (ii) scaling factors $\sqrt{\frac{\lambda_{i}+\sigma^{2}(t)}{\lambda_{i}+\sigma^{2}(T)}}$ that emphasize leading PCs. Consequently, if class-conditional covariances lack sufficiently discriminative structures (which is indeed true as shown in section˜D.3), generated samples will appear similar regardless of class label. This lack of class-specificity aligns with prior findings [25], which shows that PCs often capture generic image variations (e.g., foreground-background), rather than class-specific patterns.
The existence of the class-specificity gap implies these models fail to fully capture the higher-order statistics of the training data: if they did, naive conditional sampling, which by construction samples from the target conditional distribution, would already produce high quality samples, and CFG would only distort the target distribution. Linear diffusion models represent an extreme case: due to the linear constraint, they can only learn the first and second-order moments (mean and covariance) of the training data, which despite being fundamental data statistics, cannot capture the rich, nonlinear dependencies necessary for realistic generation. In particular, when covariances across classes share high similarity, samples initialized from the same noise become visually alike regardless of label.
@@
We hypothesize that real-world diffusion models inherit similar limitations. Although nonlinear diffusion models surely learn beyond second-order statistics, as discussed in section˜2.1, for high to moderate noise levels, they can be well approximated by linear models, especially under limited model capacity or insufficient training. Indeed, Figure˜2 (top row) and Figure˜14, 15 demonstrate that linear models reproduce the coarse-grained structures of nonlinear diffusion samples, implying that the covariance structure plays a significant role in shaping the high-level features of the generated samples. These observations reflect a well-known simplicity bias, where deep networks favor learning low-order, linearly structured representations over complex, higher-order dependencies [28]. Hence, if the covariances are indistinct across classes, sample quality can be limited even in nonlinear models (see section˜D.3 for more discussion).
@@
As quantitatively shown in Figure˜3, CFG significantly increases the inter‑class separation: FID (colored in green) between different generated classes rises. Qualitatively, Figure˜2 (bottom row) shows that CFG substantially improves both linear and nonlinear models, producing visibly better samples with enhanced class-specific structures. Similar effects of CFG across both linear and nonlinear models motivate us to use a linear model as a simplified prototype to analyze how CFG reshapes the generation process and why it is effective.
@@ 3.2 How Linear CFG Leads to Distinct Generations
Here, $f_{c}(\bm{x}_{t},t)$ is the standard conditional score, and $g_{cpc}(\bm{x}_{t},t)$ plus $g_{mean}(t)$ form the CFG guidance (derivation for the decomposition is provided in section˜B.2). Let $\bm{V}_{\sigma(t)}\hat{\bm{\Lambda}}_{\sigma(t)}\bm{V}_{\sigma(t)}^{T}$ be the eigen decomposition of $\tilde{\bm{\Sigma}}_{c,t}-\tilde{\bm{\Sigma}}_{uc,t}$, whose spectrum contains both positive and negative eigenvalues (see Figure˜19), $g_{cpc}(\bm{x}_{t},t)$ can be split accordingly into positive and negative CPC components:
@@
where $\bm{V}_{\sigma(t),+}$ and $\bm{V}_{\sigma(t),-}$ contain eigenvectors $\bm{v}_{+,i}$ and $\bm{v}_{-.i}$ corresponding to positive and negative eigenvalues $\hat{\bm{\Lambda}}_{\sigma(t),+}$ and $\hat{\bm{\Lambda}}_{\sigma(t),-}$ respectively. As discussed in section˜2.3, $\tilde{\bm{\Sigma}}_{c,t}$ and $\tilde{\bm{\Sigma}}_{uc,t}$ are up to a scaling factor $\sigma^{2}(t)$ equivalent to the conditional and unconditional posterior covariances of $p_{\text{data}}(\bm{x}|\bm{c})=\mathcal{N}(\bm{\mu}_{c},\bm{\Sigma}_{c})$ and $p_{\text{data}}(\bm{x})=\mathcal{N}(\bm{\mu}_{uc},\bm{\Sigma}_{uc})$. Hence, $\bm{V}_{\sigma(t)}$ are the CPCs which contrast between $X\sim p_{data}(\bm{x}|\bm{x}_{t},\bm{c})$ and $Y\sim p_{\text{data}}(\bm{x}|\bm{x}_{t})$. Specifically, $\bm{V}_{\sigma(t),+}$ captures directions of higher conditional variance (class-specific features), while $\bm{V}_{\sigma(t),-}$ captures directions of higher unconditional variance (features more prevalent in the unconditional data).
@@
Distinctive Effects of the CFG Components. Figure˜4(a) shows that for both nonlinear (EDM) and linear models, CFG significantly enhances the characteristic pattern—a person holding a fish—of the "tench" class from ImageNet [29]. Next, we isolate the roles of each CFG term by selectively enabling only one at a time within the linear model. In the following discussion, we omit the negative sign in (12) since the ODE runs backward in time:
@@
  • The positive CPC term (13) projects $\bm{x}_{t}-\bm{\mu}_{c}$ onto the subspace spanned by the positive CPCs, i.e., the eigenvectors $\bm{v}_{+,i}$ associated with positive eigenvalues, with each component scaled by its eigenvalue $\hat{\lambda}_{+,i}$ and the guidance strength $\gamma$. Since $\hat{\lambda}_{+,i}\geq 0$, (13) is added to $\bm{x}_{t}$, i.e., the components of $\bm{x}_{t}-\bm{\mu}_{c}$ that align with the positive CPCs, which represent the class-specific features, are amplified.  Figure˜4 (b) (second column) show the first 25 positive CPCs of $\bm{\Sigma}_{c}-\bm{\Sigma}_{uc}$  [^1] Although $\bm{V}_{\sigma(t)}$ depends on $\sigma(t)$, over a wide range of noise levels (especially high ones), it remains close to the eigenvectors of $\bm{\Sigma}_{c}-\bm{\Sigma}_{uc}$. We provide its full evolution across time in Figure 20. and the resulting samples. Compared to the conditional PCs of the dataset, the positive CPCs better capture the unique patterns of the class, which emerge visibly in the generated images.
@@
  • Similarly, the negative CPC term (14) projects $\bm{x}_{t}-\bm{\mu}_{c}$ onto the negative CPC directions $\bm{v}_{-,i}$. Since $\hat{\lambda}_{-,i}<0$, these components are subtracted from $\bm{x}_{t}$, suppressing features associated more strongly with the unconditional data. Figure˜4 (b) (third column) shows the first 25 negative CPCs and the resulting generations. Although visually less interpretable, these directions represent common but target-class-irrelevant features in the unconditional data. Suppressing them reduces background clutter and irrelevant content, making class-relevant structures more salient.
@@
  • In the context of linear diffusion model, it can be shown that (see proof in section˜B.2):
@@
    Figure˜4(b), fourth column, shows that the mean-shift term enhances the structure of class mean in the generated samples. However, unlike the positive CPC term, $g_{mean}(t)$ is independent of $\bm{x}_{t}$, thus producing more homogeneous samples with reduced diversity.
Figure 4: Distinct effects of different CFG components. (a) CFG substantially enhances class-specific features (in both EDM and linear diffusion). (b) Top row: PCs, positive/negative CPCs, and $\bm{\mu}_{c}-\bm{\mu}_{uc}$. Bottom row: generated samples when each component is applied in isolation. (c) One‑dimensional densities of generated samples after projection onto key directions. The left column corresponds to the linear diffusion model, whereas the right column corresponds to the EDM model. Top row: project onto leading positive CPC. Middle row: project onto negative CPC. Third row: project onto the mean-shift direction. Here we only plot the resulting histograms for the first positive and negative CPCs but the same patterns hold for subsequent CPCs. For experimental details and more results, please refer to section˜E.3.
@@
The proof is postponed to appendix˜B. Compared to the solution of the standard conditional sampling (8), the CFG guidance introduces the following two effects:
@@
In Section˜B.3, we provide an empirical validation of Theorem˜1 using a 2D synthetic dataset that satisfies Assumption 1. Here, we further verify the CFG’s effects of enhancing (suppressing) CPC components and shifting samples towards the mean-shift direction in natural image dataset through the following experiment:
@@
The above experiments are conducted on both linear and nonlinear (EDM) diffusion models. The resulting univariate distributions are shown in Figure˜4(c). Compared with naive conditional sampling, CFG shifts probability mass toward higher projection values along the positive-CPC and mean-shift directions, and toward lower values along the negative-CPC direction, indicating that the first two are amplified whereas the third is suppressed.